### v1

@@ Abstract
Multi-bit spiking neural networks (SNNs) have recently become a heated research spot, pursuing energy-efficient and high-accurate AI. However, with more bits involved, the associated memory and computation demands escalate to the point where the performance improvements become disproportionate. Based on the insight that different layers demonstrate different importance and extra bits could be wasted and interfering, this paper presents an adaptive bit allocation strategy for direct-trained SNNs, achieving fine-grained layer-wise allocation of memory and computation resources. Thus, SNN’s efficiency and accuracy can be improved. Specifically, we parametrize the temporal lengths and the bit widths of weights and spikes, and make them learnable and controllable through gradients. To address the challenges caused by changeable bit widths and temporal lengths, we propose the refined spiking neuron, which can handle different temporal lengths, enable the derivation of gradients for temporal lengths, and suit spike quantization better. In addition, we theoretically formulate the step-size mismatch problem of learnable bit widths, which may incur severe quantization errors to SNN, and accordingly propose the step-size renewal mechanism to alleviate this issue. Experiments on various datasets, including the static CIFAR and ImageNet and the dynamic CIFAR-DVS and DVS-GESTURE, demonstrate that our methods can reduce the overall memory and computation cost while achieving higher accuracy. Particularly, our SEWResNet-34 can achieve a 2.69% accuracy gain and 4.16$\times$ lower bit budgets over the advanced baseline work on ImageNet. This work will be fully open-sourced.
@@ 4.2 Formulation of the refined parametric spiking neuron
Thirdly, the spike amplitude coefficient $\alpha$ is usually defined differently among prior arts [36, 16, 43]. Here, we keep consistent with quantization’s perspective, and directly unify $V_{th}$ and $\alpha$ as identical, layer-wise, learnable quantization step size, as shown in Equations 12- 14.
@@ 4.3.2 Loss function
$\begin{split}&L_{total}=L_{task}+\lambda_{1}L(\bar{B}_{w},B^{tar}_{w})+\lambda_{2}L(\bar{T},T^{tar})\\ &+\lambda_{3}L(\bar{B}_{s},B^{tar}_{s}),\text{ where }L(x,y)=||x-y||_{2}^{2}.\end{split}$ (21)
@@ 4.4.2 Step-size renewal mechanism
Initialize the bit-width observer: The running maximum $V_{r\_{max}}$ is set to $\infty$- ∞ and the minimum $V_{r\_{min}}$ is set to $\infty$. The recorded bit width $B_{l}^{\prime}$ is set to 00.
@@
10:   $V^{\prime}_{max}$, $V^{\prime}_{min}$leftarrow Input $(X,V_{min},V_{max})$ to Algorithm 2.
@@
0:  Data to be quantized $X$, the $V_{max}$ and $V_{min}$ of $X$, the iteration number $K$, the minimum quantization integer $q_{min}$, the maximum quantization integer $q_{max}$, and the power constant $pow$.
@@
1:  $Score\leftarrow 0$, $R\leftarrow V_{max}-V_{min}$, $V^{\prime}_{max}\leftarrow V_{max}$, and $V^{\prime}_{min}\leftarrow V_{min}$.
@@
Method Architecture CIFAR 10 CIFAR 100 W/S/T Bit Budget S-ACE (G) Size (MB) Top-1 W/S/T Bit Budget S-ACE (G) Size (MB) Top-1 Dspike [26] ResNet18 16/1/2 32 17.72 22.42 93.13 16/1/2 32 17.72 22.42 71.68 16/1/4 64 35.44 22.42 93.66 16/1/4 64 35.44 22.42 73.35 MLF [14] ResNet20 16/1/4 64 165.58 34.55 94.25 16/1/4 64 165.58 34.55 - tdBN [50] ResNet19 16/1/2 32 73.28 25.25 92.34 16/1/2 32 73.28 25.25 - 16/1/4 64 146.56 25.25 92.92 16/1/4 64 146.56 25.25 - TET [9] ResNet19 16/1/2 32 73.28 25.25 94.16 16/1/2 32 73.28 25.25 72.87 16/1/4 64 165.12 25.25 94.44 16/1/4 64 165.12 25.25 74.47 Ternary Spike [16] ResNet19 16/2/1 32 330.24 25.25 95.58 16/2/1 32 330.24 25.25 78.45 ResNet20 16/2/2 64 165.58 34.55 94.48 16/2/2 64 165.58 34.55 73.41 16/2/4 128 330.24 34.55 94.96 16/2/4 128 330.24 34.55 74.02 Multi-bit Spike [43] ResNet20 16/4/1 64 165.58 34.55 94.59 16/4/1 64 165.58 34.55 75.43 16/4/4 256 660.48 34.55 94.93 16/4/4 256 660.48 34.55 76.51 16/4/6 384 990.72 34.55 95.00 16/4/6 384 990.72 34.55 - U-Quant.$\dagger$ ResNet20 4/4/1 16 41.40 8.64 96.35 4/4/1 16 41.40 8.64 79.83 4/4/2 32 82.79 8.64 96.39 4/4/2 32 82.79 8.64 80.19 Ours-111 ResNet18 1.45/1.61/1.04 2.43 2.96 2.03 95.13 1.65/1.87/1.03 3.17 4.29 2.32 75.42 Ours-222 1.97/2.04/2.01 8.1 11.59 / 6.56* 2.75 96.01 2.06/2.08/2.02 8.69 11.34 / 6.86* 2.89 75.80 Ours-111 ResNet20 1.38/1.32/1.00 1.83 7.55 2.98 96.15 2.06/1.58/1.00 3.25 13.37 4.45 79.70 Ours-222 2.13/2.00/2.08 8.87 29.75 / 15.93* 4.60 96.48 2.11/1.98/2.08 8.66 32.67 / 18.83* 4.56 80.59 Ours-341 3.01/4.00/1.00 12.08 39.49 6.51 97.02 2.88/4.04/1.00 11.64 40.60 6.23 81.30 $\dagger$ denotes plain uniform quantization, following [36]. * denotes the temporal squeezing is applied in inference. "Ours-xxx" denotes the target bit width of W/S/T.
@@
Method Architecture W/S/T Bit Budget S-ACE (G) Size (MB) Top-1 Hybrid training [33] ResNet34 16/1/250 4000 13608 43.56 61.48 Spiking ResNet [20] ResNet50 16/1/350 5600 18952 51.12 72.75 tdBN [50] ResNet34 16/1/6 96 340.39 43.56 63.72 Dspike [26] ResNet34 16/1/6 96 340.39 43.56 68.19 MS-ResNet [19] ResNet34 16/1/6 96 340.39 43.56 69.42 Attention-SNN [47] ResNet34 16/1/1 16 56.73 43.56 69.15 TET [9] Sew-ResNet34 16/1/4 64 226.93 43.56 68.00 Ternary Spike [16] ResNet34 16/2/4 128 435.86 43.56 70.74 PLIF+SEW [13] SEW-ResNet34 16/1/4 64 226.93 43.56 67.04 SEW-ResNet50 16/1/4 64 215.62 51.12 67.78 Quant. SNN [36] SEW-ResNet34 8/8/1 64 226.93 21.78 70.13 Ours-221 SEW-ResNet34 2.58/2.18/1.0 5.61 30.74 7.03 70.57 Ours-241 2.62/3.96/1.0 10.39 54.69 7.14 72.02 Ours-441 3.93/3.92/1.0 15.4 77.78 10.69 72.82
@@ 5 Experiments
In this section, we first compare our methods with the existing works using ResNet on CIFAR [24] and ImageNet-1k [7], then conduct ablation studies to verify different aspects of our methods, including (1) each proposed technology’s verification, (2) our methods’ extension to spiking self-attention [51], (3) mitigating the "temporal constraint" of the prior arts [36] on the event dataset CIFAR10-DVS [25] and DVS-GESTURE [2], (4) the learned bit allocation’s visualization, and (5) comparing the mixed-precision ANNs. All implementation details are attached in the appendix.
@@ 5.1.2 ImageNet
Target $\tau=2$ Flooring U-quant. $\tau=2$ Flooring w/o renewal $\tau=2$ Flooring w/ renewal $\tau=2$ Rounding w/ renewal $\tau=1$ Rounding w/ renewal W/S/T Top-1 W/S/T Top-1 W/S/T Top-1 W/S/T Top-1 W/S/T Top-1 1/1/1 74.82 1.80/2.14/1.0 95.12 1.80/2.09/1.0 95.58 1.46/1.34/1.0 95.71 1.38/1.32/1.0 96.15 1/2/1 10.02 1.77/2.2/1.0 95.21 1.58/2.13/1.0 96.14 1.40/2.05/1.0 96.48 1.42/2.08/1.0 96.58 2/2/1 0.1 2.10/2.07/1.0 95.50 2.09/1.97/1.0 96.04 1.88/1.95/1.0 96.45 2.08/2.08/1.0 96.51 2/3/1 89.06 2.05/3.01/1.0 96.51 2.13/3.16/1.0 96.38 1.96/3.04/1.0 96.72 2.24/3.06/1.0 96.63 3/3/1 0.1 2.98/3.02/1.0 96.18 3.01/3.07/1.0 96.43 3.03/3.01/1.0 96.82 2.89/2.87/1.0 96.90 3/4/1 96.31 3.06/3.98/1.0 96.64 2.99/4.02/1.0 96.80 2.95/4.06/1.0 97.00 3.01/4.0/1.0 97.02
@@ 5.2 Ablation study
We first continue to use ResNet20 and CIFAR10 to conduct ablation studies on the proposed techniques. Then, we migrate the proposed method to a different architecture, Spikformer [51], to further verify effectiveness. On the extension to the event dataset, we leverage our methods to mitigate the "temporal constraint" issue, which troubles the prior art [36]. Furthermore, we provide visualizations of the learned bit allocation as empirical evidence and statistically analyze the results. Finally, we compare our bit-adaptive SNNs with the ANN baselines to show the superiority of SNNs in scenarios of ultra-low-energy computation. Notably, since our work targets building efficient and accurate SNNs, we mainly focus on the effectiveness of the proposed approaches under the low bit budget settings and conduct ablation studies in this background.
@@ 5.2.2 Importance of the refined spiking neuron
5.2.3 Validation on spiking self-attention architectures
Migrating our methods to Spikformer is non-trivial because the self-attention operation is a different computing paradigm where a new quantizable tensor $atten$ (see Figure 2(a)) is produced. We would like to leave room for future work. Here, we naively apply our methods to Spikformer as depicted in Figure 2 only to validate the migratability and effecitveness of the proposed methods, and avoid competing with those state-of-the-art works that are based on sophisticated spiking self-attention architectures. We adopt the default Spikformer-4-384 [51] on CIFAR10. In Table 5, our bit-adaptive model surpasses the original Spikformer, while plain uniform quantization fails. Besides, adding the renewal mechanism further lifts accuracy.
5.2.4 Mitigating the temporal constraint
@@ A.2.1 Proof of the flooring operation incurring more quantization errors
[…]
@@
We further discuss ##\## as follows:
@@
    If $x>0$ and $\frac{x}{s}\notin\{\mathbb{Z},\mathbb{Z}-\frac{1}{2}\}$, then $[\frac{2x}{s}-(\left\lfloor\frac{x}{s}\right\rfloor+\left\lfloor\frac{x}{s}\right\rceil)]>0$ and $[\left\lfloor\frac{x}{s}\right\rceil-\left\lfloor\frac{x}{s}\right\rfloor]>0$. Thus, #$\#>0$.
@@
    If $x<0$ and $\frac{x}{s}\notin\{\mathbb{Z},\mathbb{Z}-\frac{1}{2}\}$, then $[\frac{2x}{s}-(\left\lfloor\frac{x}{s}\right\rfloor+\left\lfloor\frac{x}{s}\right\rceil)]<0$ and $[\left\lfloor\frac{x}{s}\right\rceil-\left\lfloor\frac{x}{s}\right\rfloor]<0$. Thus, #$\#>0$.
@@
    $\frac{x}{s}\in\{\mathbb{Z},\mathbb{Z}-\frac{1}{2}\}$, then $[\frac{2x}{s}-(\left\lfloor\frac{x}{s}\right\rfloor+\left\lfloor\frac{x}{s}\right\rceil)]=0$ and $[\left\lfloor\frac{x}{s}\right\rceil-\left\lfloor\frac{x}{s}\right\rfloor]>0$. Thus, #$\#=0$.
Therefore, if $x\in\{\mathbb{Z},\mathbb{Z}-\frac{1}{2}\}$, #$\#=0$. Otherwise, #$\#>0$. Plus, $P(x)>0$. We can get

### v2

@@ Abstract
[…]
@@ 3.1 Multi-bit spiking neuron
To measure the computational consumption of fully ideal neuromorphic hardware. Shen et al. [39] also propose the NS-ACE metric, which is calculated by:
$\begin{split}\text{NS-ACE}&=\sum_{w\in W,s\in S}fr_{s}\cdot n_{w,s}\cdot\text{BB}\\ &\approx\bar{fr}_{s}\cdot\sum_{w\in W,s\in S}n_{w,s}\cdot\text{BB}\\ &=\bar{fr}_{s}\cdot\text{S-ACE}.\end{split}$ (9)
Here, $fr_{s}$ ($\bar{fr}_{s}$ ) is the (average) firing rate of neurons, which is always less than $1$.
@@ 4.2 Formulation of the refined parametric spiking neuron
Thirdly, the spike amplitude coefficient $\alpha$ is usually defined differently among prior arts [39, 18, 48]. Here, we keep consistent with quantization’s perspective, and directly unify $V_{th}$ and $\alpha$ as identical, layer-wise, learnable quantization step size, as shown in Equations 13- 15.
@@ 4.3.2 Loss function
[…]
@@ 4.4.2 Step-size renewal mechanism
Initialize the bit-width observer: The running maximum $V_{r\_{max}}$ is set to -$\infty$ and the minimum $V_{r\_{min}}$ is set to $\infty$. The recorded bit width $B_{l}^{\prime}$ is set to 0.
@@
10:  $V^{\prime}_{max}$, $V^{\prime}_{min}$ $\leftarrow$ Input $(X,V_{min},V_{max})$ to Algorithm 2.
@@
0: Data to be quantized $X$, the $V_{max}$ and $V_{min}$ of $X$, the iteration number $K$, the minimum quantization integer $q_{min}$, the maximum quantization integer $q_{max}$, and the power constant powpow.
@@
[…]
@@
[…]
@@
Method Architecture W/S/T Bit Budget S-ACE (G) NS-ACE(G) Size (MB) Top-1 Hybrid training [36] ResNet34 16/1/250 4000 13608 - 43.56 61.48 Spiking ResNet [23] ResNet50 16/1/350 5600 18952 - 51.12 72.75 tdBN [55] ResNet34 16/1/6 96 340.39 - 43.56 63.72 Dspike [29] ResNet34 16/1/6 96 340.39 - 43.56 68.19 MS-ResNet [22] ResNet34 16/1/6 96 340.39 - 43.56 69.42 Attention-SNN [52] ResNet34 16/1/1 16 56.73 - 43.56 69.15 TET [11] Sew-ResNet34 16/1/4 64 226.93 - 43.56 68.00 Ternary Spike [18] ResNet34 16/2/4 128 435.86 - 43.56 70.74 PLIF+SEW [15] SEW-ResNet34 16/1/4 64 226.93 63.14 43.56 67.04 SEW-ResNet50 16/1/4 64 215.62 69.03 51.12 67.78 Quant. SNN [39] SEW-ResNet34 8/8/1 64 226.93 56.42 21.78 70.13 Ours-221 SEW-ResNet34 2.58/2.18/1.0 5.61 30.74 7.77 7.03 70.57 Ours-241 2.62/3.96/1.0 10.39 54.69 16.99 7.14 72.02 Ours-441 3.93/3.92/1.0 15.4 77.78 24.30 10.69 72.82
@@ 5 Experiments
In this section, we first compare our methods with the existing works using ResNet on CIFAR [27] and ImageNet-1k [9], then conduct ablation studies to verify different aspects of our methods, including (1) each proposed technology’s verification, (2) our methods’ extension to spiking self-attention [56], (3) mitigating the "temporal constraint" of the prior arts [39] on the event dataset CIFAR10-DVS [28] and DVS-GESTURE [2], (4) extending to the spiking audio task SHD [8], (5) the learned bit allocation’s visualization, and (6) comparing the mixed-precision ANNs. All implementation details are attached in the appendix.
@@ 5.1.2 ImageNet
Target $\tau=2$ & Flooring U-quant. $\tau=2$ & Flooring w/o renewal $\tau=2$ & Flooring w/ renewal $\tau=2$ & Rounding w/ renewal $\tau=1$ & Rounding w/ renewal W/S/T Top-1 W/S/T Top-1 W/S/T Top-1 W/S/T Top-1 W/S/T Top-1 1/1/1 74.82 1.80/2.14/1.0 95.12 1.80/2.09/1.0 95.58 1.46/1.34/1.0 95.71 1.38/1.32/1.0 96.15 1/2/1 10.02 1.77/2.2/1.0 95.21 1.58/2.13/1.0 96.14 1.40/2.05/1.0 96.48 1.42/2.08/1.0 96.58 2/2/1 0.1 2.10/2.07/1.0 95.50 2.09/1.97/1.0 96.04 1.88/1.95/1.0 96.45 2.08/2.08/1.0 96.51 2/3/1 89.06 2.05/3.01/1.0 96.51 2.13/3.16/1.0 96.38 1.96/3.04/1.0 96.72 2.24/3.06/1.0 96.63 3/3/1 0.1 2.98/3.02/1.0 96.18 3.01/3.07/1.0 96.43 3.03/3.01/1.0 96.82 2.89/2.87/1.0 96.90 3/4/1 96.31 3.06/3.98/1.0 96.64 2.99/4.02/1.0 96.80 2.95/4.06/1.0 97.00 3.01/4.0/1.0 97.02
@@ 5.2 Ablation study
We first continue to use ResNet20 and CIFAR10 to conduct ablation studies on the proposed techniques. Then, we migrate the proposed method to a different architecture, Spikformer [56], to further verify effectiveness. On the extension to the event dataset, we leverage our methods to mitigate the "temporal constraint" issue, which troubles the prior art [39]. We also extend our methods to the spiking audio task SHD [8] by incorporating the axonal delay mechanism [40, 20] Furthermore, we provide visualizations of the learned bit allocation as empirical evidence and statistically analyze the results. Finally, we compare our bit-adaptive SNNs with the ANN baselines to show the superiority of SNNs in scenarios of ultra-low-energy computation. Notably, since our work targets building efficient and accurate SNNs, we mainly focus on the effectiveness of the proposed approaches under the low bit budget settings and conduct ablation studies in this background.
@@ 5.2.1 Effectiveness of adaptive bit width and the renewal mechanism
In addition, another interesting finding is that the results of the target W/S/T=4/4/1 are better than those of the target W/S/T=4/4/2. This may imply that more temporal information contribute limited improvement or even cause over-fitting to the processing of static data when spatial information is efficiently provided and processed.
@@ 5.2.2 Importance of the refined spiking neuron
5.2.3 Validation on spiking self-attention architectures
Migrating our methods to Spikformer is non-trivial because the self-attention operation is a different computing paradigm where a new quantizable tensor attenatten (see Figure 2(a)) is produced. We would like to leave room for future work. Here, we naively apply our methods to Spikformer as depicted in Figure 2 only to validate the migratability and effecitveness of the proposed methods, and avoid competing with those state-of-the-art works that are based on sophisticated spiking self-attention architectures. We adopt the default Spikformer-4-384 [56] on CIFAR10. In Table 5, our bit-adaptive model surpasses the original Spikformer, while plain uniform quantization fails. Besides, adding the renewal mechanism further lifts accuracy.
5.2.4 Mitigating the temporal constraint
@@
Table 8: Extended experiments on the spiking audio dataset SHD.
Method Parameters (M) Top1 DMUC [41] 0.24 91.48 CNN [8] 1.01 92.40 RSNN [8] 1.97 83.20 Attention RSNN [51] 0.14 91.08 RadLIF [4] 3.9 94.62 DL256-SNN-Dloss [40] 0.21 93.55 DCLS [20] 0.21 95.07 Our DCLS-3.5/1/25 0.21 95.17 Our DCLS-4.4/1/25 0.21 95.54 "Ours DCLS-x/x/x" denotes the learned bit widths of W/S/T. Following the research tradition of spiking audio tasks, we only report the parameter numbers (M) instead of the model size (MB).
5.2.5 Extension to the spiking audio task
Instead of focusing solely on the event image datasets, including CIFAR10-DVS and DVS-GESTURE, we also extend our methods to the spiking audio dataset SHD [8]. However, directly applying SNNs to the spiking audio dataset will cause significant performance loss [40]. Prior arts have proposed that the axonal delay mechanism can solve this issue [40, 20]. Therefore, we utilize the dilated convolution with learnable spacings (DCLS) [20] to implement the axonal delay. Instead of directly adopting the local spike connection, which may require some modifications on the spiking neuron formulation [40], to realize axonal delays, DCLS utilizes the temporal-wise 1D convolution to implement the axonal delay mechanism. This means DCLS only operates on spiking feature maps without introducing any modification to our refined spiking neurons.
As shown in Table 8, our method not only outperforms the DCLS baseline, but also achieves the state-of-the-art (SoTA) model accuracy on the SHD dataset. This result shows our bit-adaptive SNNs are compatible with the axonal delay mechanism, and can significantly improve the SNN performance when handling the spiking audio data.
@@ 5.2.7 Additional comparisons with mixed-precision ANNs
Additionally, our SNN has adopted the proposed renewal mechanism and a bit different gradient calculations (see Section 4.3.1). Therefore, Table 9 can also imply the effectiveness of our proposed methods as a whole.
@@ A.2.1 Proof of the flooring operation incurring more quantization errors
$\begin{split}E(Err^{f})&=s^{2}[\sum_{j=0}^{2^{b}-3}\int_{-2^{b-1}+1+j}^{-2^{b-1}+2+j}(\frac{x}{s}-\left\lfloor\frac{x}{s}\right\rfloor)^{2}\cdot P(x)d\frac{x}{s}]\\ &+s^{2}\int_{-\infty}^{-2^{b-1}+1}[\frac{x}{s}-(-2^{b-1}+1)]^{2}\cdot P(x)d\frac{x}{s}\\ &+s^{2}\int_{2^{b-1}-1}^{+\infty}[\frac{x}{s}-(2^{b-1}-1)]^{2}\cdot P(x)d\frac{x}{s},\\ \text{where }P(x)&=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{x^{2}}{2\sigma^{2}}}\text{is the PDF of }x.\end{split}$ (A.1)
@@
We further discuss #\# as follows:
@@
    If $x>0$ and $\frac{x}{s}\notin\{\mathbb{Z},\mathbb{Z}-\frac{1}{2}\}$, then $\frac{2x}{s}-(\left\lfloor\frac{x}{s}\right\rfloor+\left\lfloor\frac{x}{s}\right\rceil)]>0$ and $\left\lfloor\frac{x}{s}\right\rceil-\left\lfloor\frac{x}{s}\right\rfloor]>0$. Thus, #>0\#>0.
@@
    If $x<0$ and $\frac{x}{s}\notin\{\mathbb{Z},\mathbb{Z}-\frac{1}{2}\}$, then $\frac{2x}{s}-(\left\lfloor\frac{x}{s}\right\rfloor+\left\lfloor\frac{x}{s}\right\rceil)]<0$ and $\left\lfloor\frac{x}{s}\right\rceil-\left\lfloor\frac{x}{s}\right\rfloor]<0$. Thus, #>0\#>0.
@@
    $\frac{x}{s}\in\{\mathbb{Z},\mathbb{Z}-\frac{1}{2}\}$, then $\frac{2x}{s}-(\left\lfloor\frac{x}{s}\right\rfloor+\left\lfloor\frac{x}{s}\right\rceil)]=0$ and $\left\lfloor\frac{x}{s}\right\rceil-\left\lfloor\frac{x}{s}\right\rfloor]>0$. Thus, #=0\#=0.
Therefore, if $x\in\{\mathbb{Z},\mathbb{Z}-\frac{1}{2}\}$, #=0\#=0. Otherwise, #>0\#>0. Plus, $P(x)>0$. We can get